Harry is 4 years older then Marvin. In six years Marvin will be 5/6 as old as Harry. How old is each boy?
Answer by ptaylor(2198)

Let x=Marvin's age now (x+6=Marvin's age six years from now)
Then x+4=Harry's age now (x+4+6=x+10=Harry's age six years from now)
So, our equation to solve is:
x+6=5/6(x+10) multiply each term by 6 and also get rid of parens
6x+36=5x+50 subtract 36 and also 5x from each side
6x-5x+36-36=5x-5x+50-36 collect like terms
x=14---------------------------Marvin's age now
x+4=14+4=18-------------------Harry's age now
CK
18 is 4 more that 14
and
14+6=5/6(18+6)
20=20
Hope this helps---ptaylor